# RAG-DDR: Optimizing Retrieval-Augmented Generation Using Differentiable Data Rewards

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) has proven its effectiveness in mitigating hallucinations in Large Language Models (LLMs) by retrieving knowledge from external resources. To adapt LLMs for RAG pipelines, current approaches use instruction tuning to optimize LLMs, improving their ability to utilize retrieved knowledge. This supervised fine-tuning (SFT) approach focuses on equipping LLMs to handle diverse RAG tasks using different instructions. However, it trains RAG modules to overfit training signals and overlooks the varying data preferences among agents within the RAG system. In this paper, we propose a Differentiable Data Rewards (DDR) method, which end-to-end trains RAG systems by aligning data preferences between different RAG modules. DDR works by collecting the rewards to optimize each agent with a rollout method. This method prompts agents to sample some potential responses as perturbations, evaluates the impact of these perturbations on the whole RAG system, and subsequently optimizes the agent to produce outputs that improve the performance of the RAG system. Our experiments on various knowledge-intensive tasks demonstrate that DDR significantly outperforms the SFT method, particularly for LLMs with smaller-scale parameters that depend more on the retrieved knowledge. Additionally, DDR exhibits a stronger capability to align the data preference between RAG modules. The DDR method makes generation module more effective in extracting key information from documents and mitigating conflicts between parametric memory and external knowledge. All codes are available at https://github.com/OpenMatch/RAG-DDR.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive capabilities in language understanding, reasoning, and planning capabilities across a wide range of natural language processing (NLP) tasks (Achiam et al., 2023; Touvron et al., 2023; Hu et al., 2024). However, LLMs usually produce incorrect responses due to hallucination (Ji et al., 2023; Xu et al., 2024c). To alleviate the problem, existing studies employ Retrieval Augmented Generation (RAG) (Lewis et al., 2020; Shi et al., 2023; Peng et al., 2023) to enhance the capability of LLMs and help LLMs access long-tailed knowledge and up-to-date knowledge from different data sources (Trivedi et al., 2023; He et al., 2021; Cai et al., 2019; Parvez et al., 2021). However, the conflict between retrieved knowledge and parametric memory usually misleads LLMs, challenging the effectiveness of RAG system (Li et al., 2022; Chen et al., 2023; Asai et al., 2024b).
@@
To ensure the effectiveness of RAG systems, existing research has focused on developing various agents to enhance the retrieval accuracy (Gao et al., 2024; Xu et al., 2023b; Jiang et al., 2023; Xu et al., 2024b). These approaches aim to refine retrieval results through query reformulation, reranking candidate documents, summarizing the retrieved documents or performing additional retrieval steps to find more relevant information (Yan et al., 2024; Trivedi et al., 2023; Asai et al., 2023; Yu et al., 2023a). To optimize the RAG system, the methods independently optimize different RAG modules by using the EM method (Singh et al., 2021; Sachan et al., 2021) or build the instruct tuning dataset for Supervised Fine-Tuning (SFT) these LLM based RAG modules (Lin et al., 2023; Asai et al., 2023). However, these SFT-based methods usually train LLMs to overfit the training signals and face the catastrophic forgetting problem (Luo et al., 2023b).
@@
Current research further aims to optimize these RAG modules for aligning their data preferences and primarily focuses on optimizing a two-agent framework, which consists of a retriever and a generator. Typically, these systems train only the retriever to supply more accurate documents to satisfy the data preference of the generator (Shi et al., 2023; Yu et al., 2023b). Aligning data preferences between the retriever and generator through refining retrieved knowledge is a straightforward approach to improving the effectiveness of RAG models. Nevertheless, the generator still faces the knowledge conflict, making LLMs not effectively utilize the retrieved knowledge during generation (Xie et al., 2024). Thus optimizing each RAG module by using the reward from the entire system during training, is essential for building a more tailored RAG system.
@@
This paper introduces a Differentiable Data Rewards (DDR) method for end-to-end optimizing agents in the RAG system using the DPO (Rafailov et al., 2024) method. DDR uses a rollout method (Kocsis & Szepesvári, 2006) to collect the reward from the overall system for each agent and optimizes the agent according to the reward. Specifically, we follow Asai et al. (2024a) and build a typical RAG system to evaluate the effectiveness of our DDR model. It consists of a knowledge refinement module for selecting retrieved documents and a generation module for producing responses based on queries and refined knowledge. Then we conduct the RAG-DDR model by optimizing the two-agent based RAG system using DDR. Throughout the optimization process, we use the reward from the entire RAG system and iteratively optimize both the generation and knowledge refinement modules to align data preferences across both agents.
@@
Our experiments on various Large Language Models (LLMs) demonstrate that Differentiable Data Rewards (DDR) outperforms all baseline models, achieving significant improvements over previous method (Lin et al., 2023) in a range of knowledge-intensive tasks. DDR can effectively retrofit LLMs for the RAG modeling and help generate higher-quality responses of an appropriate length. Our further analyses show that the effectiveness of our RAG-DDR model primarily derives from the generation module, which is optimized by the reward from the RAG system. Our DDR optimized generation module is more effective in capturing crucial information from retrieved documents and alleviating the knowledge conflict between external knowledge and parametric memory. Further analyses show that the effectiveness of DDR optimized RAG systems can be generalized even when additional noisy documents are incorporated during response generation.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG) is widely used in various real-world applications, such as open-domain QA (Trivedi et al., 2023), language modeling (He et al., 2021), dialogue (Cai et al., 2019), and code generation (Parvez et al., 2021). RAG models retrieve documents from external corpus (Karpukhin et al., 2020; Xiong et al., 2021) and then augment the LLM’s generation by incorporating documents as the context for generation (Ram et al., 2023) or aggregating the output probabilities from the encoding pass of each retrieved documents (Shi et al., 2023). They help LLMs alleviate hallucinations and generate more accurate and trustworthy responses (Jiang et al., 2023; Xu et al., 2023a; Luo et al., 2023a; Hu et al., 2023; Kandpal et al., 2023). However, retrieved documents inevitably incorporate noisy information, limiting the effectiveness of RAG systems in generating accurate responses (Xu et al., 2023b; 2024b; Longpre et al., 2021; Liu et al., 2024b).
@@
Some studies have demonstrated that the noise from retrieved documents can mislead LLMs, sometimes resulting in degraded performance even on some knowledge-intensive tasks  (Foulds et al., 2024; Shuster et al., 2021; Xu et al., 2024b). Such a phenomenon primarily derives from the knowledge conflict between parametric knowledge of LLMs and external knowledge (Jin et al., 2024; Longpre et al., 2021; Chen et al., 2022; Xie et al., 2024; Wu et al., 2024). Xie et al. (2024) have demonstrated that LLMs are highly receptive to external evidence when external knowledge conflicts with the parametric memory (Xie et al., 2024). Thus, lots of RAG models focus on building modular RAG pipelines to improve the quality of retrieved documents (Gao et al., 2024). Most of them aim to conduct more accurate retrieval models by employing a retrieval evaluator to trigger different knowledge refinement actions (Yan et al., 2024), prompting LLMs to summarize the query-related knowledge from retrieved documents (Yu et al., 2023a) or training LLMs to learn how to retrieve and utilize knowledge on-demand by self-reflection (Asai et al., 2023).
@@
Optimizing RAG system is a crucial research direction to help generate more accurate responses. Previous work builds a RAG system based on pretrained language models and conducts an end-to-end training method (Singh et al., 2021; Sachan et al., 2021). They regard retrieval decisions as latent variables and then iteratively optimize the retriever and generator to fit the golden answers. Recent research primarily focuses on optimizing LLMs for RAG. INFO-RAG (Xu et al., 2024a) focuses on enabling LLMs with the in-context denoising ability by designing an unsupervised pretraining method to teach LLMs to refine information from retrieved contexts. RA-DIT (Lin et al., 2023) builds a supervised training dataset and then optimizes the retriever and LLM by instruct tuning. However, these training methods focus on training LLMs to fit the training signals and face the issue of catastrophic forgetting during instruct tuning (Luo et al., 2023b).
@@
Reinforcement Learning (RL) algorithms (Schulman et al., 2017), such as Direct Preference Optimization (DPO) (Rafailov et al., 2024), are widely used to optimize LLMs for aligning with human preferences and enhancing the consistency of generated responses (Putta et al., 2024). Agent Q integrates MCTS and DPO to allow agents to learn from both successful and unsuccessful trajectories, thereby improving their performance in complex reasoning tasks (Putta et al., 2024). STEP-DPO further considers optimizing each inference step of a complex task as the fundamental unit for preference learning, which enhances the long-chain reasoning capabilities of LLMs (Lai et al., 2024). While these models primarily target the optimization of individual agents to improve response accuracy at each step, they do not focus on the effectiveness of data alignment within the multi-agent system. Instead of using SFT methods for RAG optimization (Lin et al., 2023), this paper focuses on using the DPO method to avoid overfitting the training signals and align data preferences across different agents, which is different from above RL-based optimization methods.
@@ 3 RAG Training with Differentiable Data Rewards
This section introduces the Differentiable Data Rewards (DDR) method. We first introduce the DDR method, which optimizes the agent system by aligning the data preferences between agents (Sec. 3.1). Then we utilize knowledge refinement and generation modules to build the RAG pipeline and utilize the DDR method to optimize agents in this system (Sec. 3.2).
@@ 3.1 Data Preference Learning with Differentiable Data Rewards
Data Propagation. During communication, the $t$-th agent, $V_{t}$, acts both as a sender and a receiver. Agent $V_{t}$ receives data from agent $V_{t-1}$ and simultaneously passes data to agent $V_{t+1}$:
@@
where $V_{t}\xrightarrow{y_{t}}V_{t+1}$ denotes that the agent generates one response $y_{t}$ of the maximum prediction probability and sends it to the agent $V_{t+1}$. $x\rightsquigarrow$ and $\rightsquigarrow y_{T}$ represent sending the input $x$ to the agent system $\mathcal{V}$ and getting the output $y_{T}$ from $\mathcal{V}$. The performance of the agent system $\mathcal{V}$ can be evaluated by calculating the quality score $S(y_{T})$ of the final output $V_{T}$ of the agent system $\mathcal{V}$.
Differentiable Data Reward. Unlike conventional supervised fine-tuning approaches (Lin et al., 2023), DDR tries to optimize the agent $V_{t}$ to align the data preference $V_{t+1:T}$, making the agent system produce a better response with higher evaluation score $S(y_{T})$ of the whole RAG system.
To optimize the agent $V_{t}$, DDR aims to propagate the system reward to train the targeted agent. Specifically, we first instruct $V_{t}$ to sample multiple outputs $\tilde{y}_{t}$, which incorporate some perturbations into the agent system. Then we calculate the reward $r(x,\tilde{y}_{t})$ with a rollout process (Kocsis & Szepesvári, 2006). In detail, we regard the agents $\mathcal{V}_{t+1:T}$ as the evaluation model, feed $\tilde{y}_{t}$ to this subsystem $\mathcal{V}_{t+1:T}$ and calculate the evaluation score $S(y_{T})$ of the generated output $y_{T}$:
@@
where $\sigma$ is the Sigmoid function. The parameters of $V_{t}$ can be trained using the DPO (Rafailov et al., 2024) training loss, aiding the DDR in identifying optimization directions by contrastively learning from the positive ($\tilde{y}_{t}^{+}$) and negative ($\tilde{y}_{t}^{-}$) outputs:
@@ 3.2 Optimizing a Specific RAG System through DDR
Given a query $q$ and a set of retrieved documents $D=\{d_{1},\dots,d_{n}\}$, we build a RAG system by employing a knowledge refinement module ($V_{\text{KR}}$) to filter unrelated documents and a generation module ($V_{\text{Gen}}$) to produce a response. These two modules can be represented by a two-agent system:
@@
The response that achieves the highest evaluation score $S(\tilde{y}_{\text{Gen}})$ is considered positive ($\tilde{y}_{\text{Gen}}^{+}$), while the lowest evaluation score is considered negative ($\tilde{y}_{\text{Gen}}^{-}$). The generation module $V_{\text{Gen}}$ is optimized to maximize the probability of generating the positive response $P(\tilde{y}_{\text{Gen}}^{+}>\tilde{y}_{\text{Gen}}^{-}|q,\tilde{D})$ to win a higher reward. By generating responses $\tilde{y}_{\text{Gen}}$ based on documents or the query alone, LLMs can learn to balance internal and external knowledge, alleviating the problems related to knowledge conflicts.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) has proven its effectiveness in mitigating hallucinations in Large Language Models (LLMs) by retrieving knowledge from external resources. To adapt LLMs for the RAG systems, current approaches use instruction tuning to optimize LLMs, improving their ability to utilize retrieved knowledge. This supervised fine-tuning (SFT) approach focuses on equipping LLMs to handle diverse RAG tasks using different instructions. However, it trains RAG modules to overfit training signals and overlooks the varying data preferences among agents within the RAG system. In this paper, we propose a Differentiable Data Rewards (DDR) method, which end-to-end trains RAG systems by aligning data preferences between different RAG modules. DDR works by collecting the rewards to optimize each agent in the RAG system with the rollout method, which prompts agents to sample some potential responses as perturbations, evaluates the impact of these perturbations on the whole RAG system, and subsequently optimizes the agent to produce outputs that improve the performance of the RAG system. Our experiments on various knowledge-intensive tasks demonstrate that DDR significantly outperforms the SFT method, particularly for LLMs with smaller-scale parameters that depend more on the retrieved knowledge. Additionally, DDR exhibits a stronger capability to align the data preference between RAG modules. The DDR method makes the generation module more effective in extracting key information from documents and mitigating conflicts between parametric memory and external knowledge. All codes are available at https://github.com/OpenMatch/RAG-DDR.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive capabilities in language understanding, reasoning, and planning across a wide range of natural language processing (NLP) tasks (Achiam et al., 2023; Touvron et al., 2023; Hu et al., 2024). However, LLMs usually generate incorrect responses due to hallucination (Ji et al., 2023; Xu et al., 2024c). To alleviate this problem, existing studies employ Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Shi et al., 2023; Peng et al., 2023) to enhance the ability of LLMs and help LLMs access long-tailed knowledge and up-to-date knowledge from different data sources (Trivedi et al., 2023; He et al., 2021; Cai et al., 2019; Parvez et al., 2021). However, the conflict between retrieved knowledge and parametric memory usually misleads LLMs, challenging the effectiveness of the RAG system (Li et al., 2022; Chen et al., 2023; Asai et al., 2024).
@@
To ensure the effectiveness of RAG systems, existing research has focused on developing various agents to obtain high-quality external knowledge (Gao et al., 2024; Xu et al., 2024a; Jiang et al., 2023; Xu et al., 2024b), which can refine retrieval documents through query reformulation, reranking candidate documents, summarizing the retrieved documents, or performing additional actions to obtain more relevant information for LLMs (Yan et al., 2024; Trivedi et al., 2023; Asai et al., 2023; Yu et al., 2023a). To further optimize the RAG system, some methods independently optimize different RAG modules by using the EM method (Singh et al., 2021; Sachan et al., 2021) or build the instruction tuning dataset for Supervised Fine-Tuning (SFT) these LLMs-based RAG modules (Lin et al., 2023; Asai et al., 2023). However, these SFT-based methods usually make LLMs overfit the training signals and face the catastrophic forgetting problem (Luo et al., 2023b).
@@
Furthermore, current research aims to optimize RAG modules by aligning their data preferences and primarily focuses on optimizing a two-agent RAG system, which consists of retrieval and generation modules. Typically, these methods train only the retrieval module by using the preference signals from the generation module, making the retrieval module supply more accurate documents to satisfy the data preference of the generation module (Shi et al., 2023; Yu et al., 2023b). However, these methods only optimize the retrieval module and overlook that the generation module still faces the knowledge conflict (Xie et al., 2024), making the generation module’s outputs do not align with the data preferences of the RAG system. Thus, optimizing and aligning the data preferences of each module in the RAG system is essential for building a more tailored RAG system.
@@
This paper introduces a Differentiable Data Rewards (DDR) method for end-to-end optimizing each agent in the RAG system using the DPO (Rafailov et al., 2024) method. DDR uses a rollout method (Kocsis & Szepesvári, 2006) to collect the reward from the overall system for each agent and optimizes the agent according to the reward. Specifically, we follow Asai et al. (2023) and build a typical RAG system to evaluate the effectiveness of DDR model. This RAG system consists of a knowledge refinement module for selecting retrieved documents and a generation module for producing responses based on the query and refined knowledge. Then we conduct the RAG-DDR model by optimizing the two-agent based RAG system using DDR. We use the reward from the entire RAG system and iteratively optimize both the generation and knowledge refinement modules to align data preferences across both agents, enabling the RAG system to generate accurate responses.
@@
Our experiments on various Large Language Models (LLMs) demonstrate that Differentiable Data Rewards (DDR) outperforms all baseline models, achieving significant improvements over the previous method (Lin et al., 2023) in a range of knowledge-intensive tasks. DDR can effectively retrofit LLMs for the RAG modeling and help LLMs generate higher-quality responses of an appropriate length. Our further analyses show that the effectiveness of our RAG-DDR model primarily derives from the generation module, which is optimized by the reward from the RAG system. The DDR-optimized generation module is more effective in capturing crucial information from retrieved documents and alleviating the knowledge conflict between external knowledge and parametric memory. Further analyses show that the effectiveness of DDR-optimized RAG systems can be generalized even when additional noisy documents are incorporated during response generation.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG) is widely used in various real-world tasks, such as open-domain QA (Trivedi et al., 2023), language modeling (He et al., 2021), dialogue (Cai et al., 2019), and code generation (Parvez et al., 2021). RAG models retrieve documents from external corpus (Karpukhin et al., 2020; Xiong et al., 2021) and then augment the LLM’s generation by incorporating documents as the context of input (Ram et al., 2023) or aggregating the output probabilities from the encoding pass of each retrieved document (Shi et al., 2023). They help LLMs alleviate hallucinations and generate more accurate and trustworthy responses (Jiang et al., 2023; Xu et al., 2023; Luo et al., 2023a; Hu et al., 2023; Kandpal et al., 2023). However, retrieved documents inevitably incorporate noisy information, limiting the effectiveness of RAG systems in generating accurate responses (Xu et al., 2024a; b; Longpre et al., 2021; Liu et al., 2024b).
@@
Some studies have demonstrated that the noise from retrieved documents can mislead LLMs, sometimes resulting in degraded performance even on some knowledge-intensive tasks  (Foulds et al., 2024; Shuster et al., 2021; Xu et al., 2024b). This issue primarily derives from the knowledge conflict between the parametric knowledge of LLMs and external knowledge (Jin et al., 2024; Longpre et al., 2021; Chen et al., 2022; Xie et al., 2024; Wu et al., 2024). Xie et al. (2024) have demonstrated that LLMs are highly receptive to external evidence when external knowledge conflicts with the parametric memory (Xie et al., 2024). Thus, lots of RAG models focus on building modular RAG pipelines to improve the quality of retrieved documents (Gao et al., 2024). Most of them aim to conduct more accurate retrieval models by employing a retrieval evaluator to trigger different knowledge refinement actions (Yan et al., 2024), prompting LLMs to summarize the query-related knowledge from retrieved documents (Yu et al., 2023a) or training LLMs to learn how to retrieve and utilize knowledge on-demand by self-reflection (Asai et al., 2023).
@@
Optimizing the RAG system is a crucial research direction to help LLMs generate more accurate responses. Previous work builds a RAG system based on pretrained language models and conducts an end-to-end training method (Singh et al., 2021; Sachan et al., 2021). They regard retrieval decisions as latent variables and then iteratively optimize the retriever and generator to fit the golden answers. Recent research primarily focuses on optimizing the LLMs in RAG systems. INFO-RAG (Xu et al., 2024a) focuses on enabling LLMs with the in-context denoising ability by designing an unsupervised pretraining method to teach LLMs to refine information from retrieved contexts. RA-DIT (Lin et al., 2023) builds a supervised training dataset and then optimizes the retriever and LLM by instruction tuning. However, these training methods focus on training LLMs to fit the training signals and face the issue of catastrophic forgetting during instruction tuning (Luo et al., 2023b).
@@
Reinforcement Learning (RL) algorithms (Schulman et al., 2017), such as Direct Preference Optimization (DPO) (Rafailov et al., 2024), are widely used to optimize LLMs for aligning with human preferences and enhancing the consistency of generated responses (Putta et al., 2024). Agent Q integrates MCTS and DPO to allow agents to learn from both successful and unsuccessful trajectories, thereby improving their performance in complex reasoning tasks (Putta et al., 2024). STEP-DPO further considers optimizing each inference step of a complex task as the fundamental unit for preference learning, which enhances the long-chain reasoning capabilities of LLMs (Lai et al., 2024). While these models primarily target the optimization of individual agents to improve response accuracy at each step, they do not focus on the effectiveness of data alignment within the multi-agent system. Instead of using SFT methods for RAG optimization (Lin et al., 2023), this paper focuses on using the DPO method to avoid overfitting the training signals and aligning data preferences across different agents, which is different from the above RL-based optimization methods.
@@ 3 RAG Training with Differentiable Data Rewards
This section introduces the Differentiable Data Rewards (DDR) method. We first introduce the DDR method, which optimizes the RAG system by aligning the data preferences between different agents (Sec. 3.1). Then, we employ knowledge refinement and generation modules to build the RAG system and utilize the DDR method to optimize the agents in this RAG system (Sec. 3.2).
@@ 3.1 Data Preference Learning with Differentiable Data Rewards
Data Propagation. During communication, the $t$-th agent, $V_{t}$, acts as both a sender and a receiver. Agent $V_{t}$ receives data from agent $V_{t-1}$ and simultaneously passes data to agent $V_{t+1}$:
@@
where $V_{t}\xrightarrow{y_{t}}V_{t+1}$ denotes that the agent generates one response $y_{t}$ of the maximum prediction probability and sends it to the agent $V_{t+1}$. $x\rightsquigarrow$ and $\rightsquigarrow y_{T}$ represent sending the input $x$ to the RAG system $\mathcal{V}$ and getting the final output $y_{T}$ from $\mathcal{V}$. The performance of the RAG system $\mathcal{V}$ can be evaluated by calculating the quality score $S(y_{T})$ of the final output $y_{T}$.
Differentiable Data Reward. Unlike conventional supervised fine-tuning approaches (Lin et al., 2023), DDR tries to optimize the agent $V_{t}$ to align its data preference with $V_{t+1:T}$, making the RAG system produce a better response with a higher evaluation score $S(y_{T})$.
To optimize the agent $V_{t}$, DDR aims to propagate the system reward to train the targeted agent. Specifically, we first instruct $V_{t}$ to sample multiple outputs $\tilde{y}_{t}$, which incorporate some perturbations into the RAG system. Then, we calculate the reward $r(x,\tilde{y}_{t})$ with a rollout process (Kocsis & Szepesvári, 2006). In detail, we regard the agents $\mathcal{V}_{t+1:T}$ as the evaluation model, feed $\tilde{y}_{t}$ to this subsystem $\mathcal{V}_{t+1:T}$ and calculate the evaluation score $S(y_{T})$ of the final output $y_{T}$:
@@
where $\sigma$ is the Sigmoid function. The parameters of $V_{t}$ can be trained using the DPO (Rafailov et al., 2024) training loss, aiding the DDR model in identifying optimization directions by contrastively learning from the positive ($\tilde{y}_{t}^{+}$) and negative ($\tilde{y}_{t}^{-}$) outputs:
@@ 3.2 Optimizing a Specific RAG System through DDR
Given a query $q$ and a set of retrieved documents $D=\{d_{1},\dots,d_{n}\}$, we build a RAG system by employing a knowledge refinement module ($V_{\text{KR}}$) to filter unrelated documents and a generation module ($V_{\text{Gen}}$) to answer the question. These two modules can be represented by a two-agent system:
@@
The response that achieves the highest evaluation score $S(\tilde{y}_{\text{Gen}})$ is considered positive ($\tilde{y}_{\text{Gen}}^{+}$), while the lowest evaluation score is considered negative ($\tilde{y}_{\text{Gen}}^{-}$). The generation module $V_{\text{Gen}}$ is optimized to maximize the probability of generating the positive response $P(\tilde{y}_{\text{Gen}}^{+}>\tilde{y}_{\text{Gen}}^{-}|q,\tilde{D})$ to win a higher reward. By generating responses $\tilde{y}_{\text{Gen}}$ based on documents or only the query, LLMs can learn to balance internal and external knowledge, alleviating the knowledge conflict problem.